OpenStudy (anonymous):

Calculate the constant acceleration a in g’s which the catapult of an aircraft carrier must provide to produce a launch velocity of 188 mi/hr in a distance of 299 ft. Assume that the carrier is at anchor.

OpenStudy (shane_b):

First, I'd convert 188 mi/hr to ft/s. You should end up with about ~275.7 ft/s. So now write down all the values you know: Vfinal = 275.7 ft/s Vinitial = 0 ft/s distance = 299ft Rearrange the kinematic equation below to solve for "a": \[\large V_f^2=V_i^2+2ad\]\[\large a = \frac{V_f^2-V_i^2}{2d}\]Now just plug in Vf, Vi and d to solve.

OpenStudy (shane_b):

BTW: That will give you the acceleration in ft/s^2. You can convert that to "g"s by dividing it by 32 since 1 g is 32 ft/s^2.